OpenStudy (lizeth4101):

Predict the total cost if you bought 20 books. Books Bought 1 3 5 7 Total Amount Spent $ 21 39 57 75 a)$192 b)$420 c)$200 d)$310

OpenStudy (ad7):

Every two books more you have to pay $18 more. So each book is $9, except the first, which was $21. 21+19. 9 equals A

OpenStudy (camerondoherty):

Ok, sooo We start out with an initial amount of 21 right? So the starting point for us is $21 The data shows not every book, but every two books It is not every three books because it starts with 1, not 0 1 + 2 = 3 3 + 2 = 5 etc. So, from $21 to $39, how many dollars did we increase Same with 3 to 5 From $39 to $57, how many dollars did we increase?

OpenStudy (lizeth4101):

you increase $18 every two books right?

OpenStudy (camerondoherty):

Yep! So since it is every two books, you divide by 2 So that would mean how much is each book if each book is $18?

OpenStudy (lizeth4101):

$9

OpenStudy (camerondoherty):

Yep :) Now we want to find the total amount she spent SInce the first book was at an extra amount we'll set that one aside To find the amount she spent, we would need to multiply 9 x 19 The number is 19 and not 20 because we set aside that first payment of $21 So, what is 9 x 19?

OpenStudy (lizeth4101):

171

OpenStudy (camerondoherty):

Perfect c: Now you just add the beggining payment to that, which was the $21 we started with So 171 + 21 = ?

OpenStudy (lizeth4101):

192

OpenStudy (camerondoherty):

And that's your answer c: Great Job!

OpenStudy (lizeth4101):

thx!!